Four years ago, Katie was twice as old as Anne was then. In 6 years, Anne will be the same age that Katie is now. How old is each now? Thanks so much!
Answer by dabanfield(803)

Four years ago, Katie was twice as old as Anne was then. In 6 years, Anne will be the same age that Katie is now. How old is each now? Thanks so much!
Let K be Katie's current age and A be Anne's current age.
Then from the first statement we know:
1.) K-4 = 2*(A-4)
From the second statement we know:
2.) A+6 = K
Based on equation 2.) substitute A+6 for K in equation 1.):
(A+6) - 4 = 2*(A-4) or
A+2 = 2A-8
Solve the above for A and then calculate K from 2.) as K = A+6